OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points? (–5.5, 6.1), (–2.5, 3.1) (1 point)

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

(6.1-3.1)/(-5.5--2.5)= 3/(-3)= -1